OpenStudy (kj4uts):

If h(x)=(fog)(x) and h(x)=^3√x+3, find g(x) if f(x)=^3√x+2.

OpenStudy (kj4uts):

Problem attached below. Please help and explain. Thank you to those who help!

OpenStudy (anonymous):

\[h(x) = \sqrt[3]{x+3};f(x) = \sqrt[3]{x+2}\] \[fog(x) = f(g(x)) = \sqrt[3]{g(x) +2}\] \[f(g(x)) = \sqrt[3]{g(x) + 2} = h(x) = \sqrt[3]{x+3}\] Cube both sides: \[g(x) +2 =x + 3\] \[g(x) = x + 3 - 2 = x +1\] \[g(x) = x +1\]

OpenStudy (kj4uts):

im still confused but thank you.

OpenStudy (anonymous):

Which part are you confused with?